# Near-Optimal Decision Trees in a SPLIT Second

### v2

@@ 1 Introduction
Decision tree optimization is core to interpretable machine learning (Rudin et al., 2022). Simple decision trees present the entire model reasoning process transparently, directly allowing faithful interpretations of the model (Arrieta et al., 2020). This helps users choose whether to trust the model and to critically examine any perceived flaws.
@@
Figure 1: An illustration of the power of our optimization algorithm. We train $3$ decision trees on the Bike dataset, with the aim of predicting bike rentals in Washington DC in a given time period. A greedy tree is fast but suboptimal. An optimal tree is well performing but very slow. Our algorithm strikes the perfect balance, providing well performing trees in a SPLIT second, orders of magnitude faster than optimal approaches seen in literature.
@@
Optimizing the performance of decision trees while preserving their simplicity presents a significant challenge. Traditional greedy methods scale linearly with both dataset size and the number of features (Breiman, 1984; Quinlan, 2014).
@@
However, these methods tend to yield suboptimal results, lacking general guarantees on either sparsity or accuracy. Recent advances in decision tree algorithms use dynamic programming techniques combined with branch-and-bound strategies, offering solutions that are faster than brute-force approaches and provably optimal (Lin et al., 2020; Aglin et al., 2020; Demirović et al., 2022; McTavish et al., 2022). In fact, Demirović et al. (2022) and van der Linden et al. (2024) reveal an average gap of $1$-$2$ percentage points between greedy and optimal trees, with Demirović et al. (2022) showing that some datasets can exhibit gaps as large as $10$ percentage points. These algorithms struggle to scale to datasets with hundreds or thousands of features or to deeper trees. It seems that we should return to greedy methods for larger-scale problems, but this would come at a loss of performance. Ideally, we should leverage greed only when it does not significantly deviate from optimality and use dynamic programming otherwise. Dynamic programming approaches build trees recursively, downward from the root. Problems farther from the root contain fewer samples and produce fewer splits. As we show, greedy splits near the root sacrifice performance, while greedy splits near the leaves produce performance close to the optimal. This suggests that we can tolerate less precision on problems close to leaves than on problems closer to the root – and that full optimization on those problems closer to the leaves yields only marginal returns relative to greedy, since we only have a few splits remaining. This has enormous implications, since the number of candidate trees increases exponentially with increases in depth; using greedy splitting closer to the leaves of the tree massively reduces the search space.
We leverage this observation to construct SPLIT (SParse Lookahead for Interpretable Trees), a family of decision tree algorithms that are over 100$\times$ faster than state of the art optimal decision tree algorithms, with negligible sacrifice in performance. They can also be tuned to a user-defined level of sparsity. Instead of searching through the entire space of decision trees up to a given depth, our algorithm performs dynamic programming with branch and bound up to only a shallow “lookahead” depth, conditioned on all splits henceforth being chosen greedily.
@@
- We develop a family of decision tree algorithms that scale with the dataset size and number of features comparably to standard greedy algorithms but produce trees that are as accurate and sparse as optimal ones (e.g., Lin et al., 2020).
- We extend our decision tree algorithms to allow scalable, accurate approximations of the Rashomon set of decision trees (Breiman, 2001; Xin et al., 2022).
@@ 2 Related Work
Consistent with recommendations from Rudin et al. (2022); Costa & Pedreira (2023), we emphasize sparsity, expressed in terms of the number of leaves, as the primary mechanism for tree interpretability. Sparsity has a strong correlation with user comprehension (Piltaver et al., 2016). Zhou et al. (2018) fit a regression model to user-reported interpretability for decision trees, also finding that trees with fewer leaves were more interpretable. They also found that deep, sparse trees were more interpretable than shallow trees with the same sparsity. Izza et al. (2022) provides a way to use a sparse decision tree to provide succinct individual explanations. However, finding deep, sparse trees with existing methods can be computationally infeasible. We bridge this gap – our algorithms are capable of finding sparse trees without constraining them to be shallow.
@@
A long line of work explores greedy algorithms such as CART (Breiman, 1984) and C4.5 (Quinlan, 2014). These methods first define a heuristic feature quality metric such as the Gini impurity score (Breiman, 1984) or the information gain (Quinlan, 2014) rather than choosing a global objective function. At every decision node, the feature with the highest quality is chosen as the splitting feature. This process is repeated until a termination criteria is reached. One such criteria often used is the minimum support of each leaf. Trees can then be postprocessed with pruning methods.
@@
Among the many methods for globally optimizing trees, Branch-and-bound approaches with dynamic programming are state of the art for scalability, because they exploit the structure of decision trees (Costa & Pedreira, 2023; Lin et al., 2020; Demirović et al., 2022; McTavish et al., 2022; Aglin et al., 2020). While many other methods exist for optimizing trees, such as MIP solvers (Bertsimas & Dunn, 2017; Verwer & Zhang, 2019), we focus our discussion and comparison of globally optimal decision tree methods on the currently fastest types of approaches – dynamic programming with branch and bound (DPBnB). These approaches search through the space of decision trees while tracking lower and upper bounds of the overall objective at each split to reduce the search space. They can find optimal trees on medium-sized datasets with tens of features and shallow maximum tree depths (Sullivan et al., 2024; Aglin et al., 2020; Lin et al., 2020; Demirović et al., 2022). Aglin et al. (2020) uses a DPBnB method with advanced caching techniques to find optimal decision trees, though it does not explicitly optimize for sparsity. In contrast, Lin et al. (2020); Hu et al. (2019) use a DPBnB approach to find a tree that optimizes a weighted combination of empirical risk and sparsity, defined by the number of leaves in the tree. McTavish et al. (2022) further enhances this approach by incorporating smart guessing strategies to construct tighter lower bounds for DPBnB, resulting in computational speedups. Demirović et al. (2022) extends the work of Aglin et al. (2020) by focusing on finding the optimal tree with a hard constraint on the number of permissible nodes, using advanced caching techniques and an optimized depth-2 decision tree solver. Mazumder et al. (2022) addresses continuous features by defining lower and upper bounds based on quantiles of feature distributions. However, their method is applicable only to shallow optimal trees with depth $\leq 3$, limiting its utility in scenarios with higher-order feature interactions.
@@
Some older approaches to greedy decision tree optimization consider multiple levels of splits before selecting the best split at a given iteration (Norton, 1989). That is, unlike the other greedy approaches, these approaches do not pick the split that optimizes a heuristic immediately. Instead, they pick a split that sets up the best possible heuristic value on the following split.
These approaches still focus on locally optimizing a heuristic measure that is not necessarily aligned with a global objective. By contrast, our method selects splits to directly optimize the sparse misclassification rate of the final tree. We globally optimize the search up to the specified lookahead depth, switching to heuristics only when deciding splits past our lookahead depth. In so doing, our method largely avoids the pathology noted in Murthy & Salzberg (1995), who note cases where their own lookahead approach results in a substantially worse tree than one constructed with a standard greedy approach. For our method, it is provably impossible for a fully greedy entropy-based method with the same constraints as our approach to achieve a better training set objective than our approach. (See Theorem A.1)
@@
Several other approaches are compatible with branch and bound techniques. Blanc et al. (2024) seek to bridge the gap between greedy and optimal decision trees by selecting a fixed subset of the top $k$ feature splits for each sub-problem. However, this framework does not explicitly account for sparsity. Further, the method is limited by using a global setting for search precision: the approach considers the same number of candidate splits at each subproblem. As we show in our experiments, there is merit to tailoring the level of search precision to parts of the search space where it is most needed. The Blossom algorithm (Demirović et al., 2023) traverses a branch and bound dependency graph structure while using greedy heuristics to guide the search order. Relative to our approach, this algorithm optimizes from the bottom up, starting with greedy splits at each level, then optimizing the splits furthest from the root first. This choice guarantees eventual optimality while giving anytime behavior, but misses out on leveraging the property motivating this work – that greedy splits are most detrimental near the top of the tree. Like the approach of Blanc et al. (2024), Blossom also does not account for sparsity.
There are a few methods that use probabilistic search techniques to optimize trees. Sullivan et al. (2024) take a Bayesian approach, finding the maximum-a-posteriori tree by optimizing over an AND/OR graph, akin to the graph used in earlier branch-and-bound methods like that of Lin et al. (2020). Although their method demonstrates strong performance, their experimental results reveal that it is not responsive to sparsity-inducing hyperparameters – accordingly, we found in our experiments that the method struggles to optimize for sparsity.
Recent work by Chaouki et al. (2024) devises a Monte Carlo Tree Search algorithm using Thompson sampling to enable online, adaptive learning of sparse decision trees. We show that our method achieves superior performance and sparsity on all datasets tested.
@@ 3 Preliminaries
We now briefly discuss sparse greedy and optimal trees. We define $T_{g}(D,d,\lambda)$ to be a decision tree of depth at most $d$ trained greedily on $D$ with sparsity penalty $\lambda$. Intuitively, this sparse greedy algorithm will make a split at a node only when the gain in overall accuracy is greater than $\lambda$. Algorithm 4 in the Appendix illustrates this procedure. Modern methods such as Lin et al. (2020); McTavish et al. (2022), on the other hand, find a tree $T$ in the space of decision trees $\mathcal{T}$ that solves the following optimization problem:
@@
$=\min_{T\in\mathcal{T}}\sum_{i=1}^{N}\frac{1}{N}\Big{(}{l\big{(}T(\mathbf{x}_{i}),y_{i}\big{)}+\lambda S(T)\Big{)}}\textrm{s.t. depth$(T)$ $\leq d$}$
where $L(T,D,\lambda)$ is the regularized loss of tree $T$ on dataset $D$, $S(T)$ is the number of leaves in $T$ and $\ell(T(\mathbf{x}),y)$ is the loss incurred by $T$ in its prediction on $\mathbf{x}$ (for this paper, we set $\ell$ to be the 0-1 loss). As discussed in Section 2, the fastest contemporary methods solve this problem using a branch-and-bound approach (Costa & Pedreira, 2023; Lin et al., 2020; Demirović et al., 2022; McTavish et al., 2022).
@@
Our work is motivated by the properties of near-optimal decision trees and allows for scalable approximation of that set. Xin et al. (2022) define the Rashomon set, denoted by $\mathcal{R}(D,\lambda,\epsilon,d)$, as the collection of all trees whose objective is within $\epsilon$ of the minimum value in Equation 1. Formally:
@@
Rashomon sets can be used for a range of downstream tasks (Rudin et al., 2024); one crucial task is the measurement of variable importance over a set of near-optimal models instead of only for a single model (Donnelly et al., 2023; Fisher et al., 2019). Reliable variable importance measures in this setting rely on minimal feature selection prior to computing the Rashomon set and minimal constraints on the tree’s depth to allow high-order interactions. Our approach can be used to accelerate the computation of a Rashomon set, supporting the feasibility of these approaches.
@@
Given a depth budget $d$, branch and bound with a sparsity penalty (Lin et al., 2020; McTavish et al., 2022) finds the optimal loss $\mathcal{L}^{*}(D,d,\lambda)$ that minimizes Equation 1.
@@
Our algorithm will assume feature vectors to be binary, i.e., $\mathbf{x}_{i}\in\{0,1\}^{K}$. Real-world datasets often have features that require discretization to fit our setting. While some methods preserve optimality (e.g., splitting at the mean between unique values in the training set), others such as bucketization (described and proven to be suboptimal in Lin et al., 2020), binning into quantiles, and feature engineering reduce the search space at the cost of optimality. In our experiments, we use threshold guessing (McTavish et al., 2022), which sacrifices optimality with respect to a real-valued dataset but maintains theoretical and empirical guarantees relative to a reference decision tree ensemble.
@@ 4 Algorithm Motivation
A key motivating property of SPLIT is that we can find high quality trees even when splitting greedily far from the root of the tree. To support this intuition, we empirically investigate how frequently near optimal trees behave greedily far from the root. To do so, we first generate the Rashomon set of decision trees for various values of sparsity penalty $\lambda$ and Rashomon bound $\epsilon$. Let $T\in\mathcal{R}(D,\lambda,\epsilon,d)$ be a tree in the Rashomon set, and let $n\in T$ be any node in $T$. Then, we compute the fraction of all nodes at a given level $\ell\leq d$ (where level 00 corresponds to the root) that were greedy (by which we mean that the split at this node in the tree is optimal with respect to information gain). This corresponds to the following proportion:
@@
Figure 2: A heatmap of the proportion of splits of trees in the Rashomon set that are greedy, stratified by level, for different ($\lambda,\epsilon$) combinations. Only $4$ levels are shown as the $5^{th}$ level corresponds to the leaf. The greyed out regions in the bottom right of a plot represent ($\lambda,\epsilon$) for which the Rashomon set did not contain any trees of that depth. Generally, as we approach the leaves, the proportion of splits appearing in $\epsilon$-optimal trees become increasingly greedy. This is especially noticeable for the Netherlands, Covertype, and COMPAS datasets.
@@ 5.1 SParse Lookahead for Interpretable Trees (SPLIT)
$\begin{cases}\lambda+\min\Bigg{\{}\frac{|D^{-}|}{|D|},\frac{|D^{+}|}{|D|}\Bigg{\}}\ \ \hskip 44.10185pt\text{if $d^{\prime}=0$}\\ \\ \begin{aligned} &\min\Bigg{\{}\lambda+\frac{|D^{-}|}{|D|},\lambda+\frac{|D^{+}|}{|D|},\\ &\min_{f\in\mathcal{F}}\Big{\{}L\Big{(}T_{g}\big{(}D(f),d^{\prime},\lambda\big{)}\Big{)}+L\Big{(}T_{g}\big{(}D(\bar{f}),d^{\prime},\lambda\big{)}\Big{)}\Big{\}}\Bigg{\}}\end{aligned}\ \ \\ \hskip 142.26378pt\text{if $d^{\prime}=d-d_{l}$}\\ \begin{aligned} &\min\Bigg{\{}\lambda+\frac{|D^{-}|}{|D|},\lambda+\frac{|D^{+}|}{|D|},\\ &\min_{f\in\mathcal{F}}\Big{\{}\mathcal{L}\Big{(}D(f),d^{\prime}-1,\lambda\Big{)}+\mathcal{L}\Big{(}D(\bar{f}),d^{\prime}-1,\lambda\Big{)}\Big{\}}\Bigg{\}}\end{aligned}\ \ \\ \hskip 142.26378pt\text{if $d^{\prime}>d-d_{l}$.}\end{cases}$ (4)
Algorithm 1 get_bounds($D$, $d_{l}$, $d$) $\to$ lb, ub
0:  $D$, $d_{l}$, $d$support, lookahead depth, current search depth}
1:  if $d=d_{l}$ then
2:     $T_{g}=$ Greedy$(D,d-d_{l},\lambda)$Find greedy tree rooted at $D$ (Alg 4 in the Appendix)}
@@
4:     $\alpha\leftarrow\frac{1}{|D|}\sum_{(x,y)\in D}\mathbf{1}[y\neq T_{g}(x)]+\lambda S(T_{g})$
@@
9:     $ub\leftarrow\lambda+\min\Big{\{}\frac{|D^{-}|}{|D|},\frac{|D^{+}|}{|D|}\Big{\}}$
@@
Algorithm 2 $\ell$, D, $\lambda$, $d_{l}$, $d$, $p$)
0:  $\ell$, $D$, $\lambda$, $d_{l}$, $d$ , $p$loss function, samples, regularizer, lookahead depth, depth budget, postprocess flag}
1:  ModifiedGOSDT = GOSDT reconfigured to use get__\__bounds (Algorithm 1) whenever it encounters a new subproblem
1:  $t_{lookahead}=$ ModifiedGOSDT$(\ell,D,\lambda,d_{l})$Call ModifiedGOSDT with depth budget $d_{l}$}
@@
We can constrain the search space to include only greedy trees past the lookahead depth by modifying the lower and upper bounds used in branch and bound (see Algorithm 1). In particular, sub-problem nodes initialized at depths up to the lookahead depth are assigned initial lower and upper bounds equivalent to that in GOSDT (Lin et al., 2020) (see Section 2). At the lookahead depth, however, the lower and upper bounds for a subproblem are fixed to be the loss of a greedy subtree trained on that subproblem. After these bound assignments, our algorithm uses the GOSDT algorithm with these new bounds to solve Equation 5.1 – this is summarized by Lines $1$-$2$ in Algorithm 2. We defer more details of the GOSDT algorithm to Section A.12 in the Appendix.
@@
Once we have solved Equation 5.1, we do not need to use greedy sub-trees past the lookahead depth. We can improve our approach by replacing these subtrees with fully optimal decision trees. Lines $3$-$9$ in Algorithm 2 illustrate this. Thus, the performance of the lookahead tree with the aforementioned greedy subtrees is just an upper bound on the objective of the tree our method ultimately finds.
Note that the renormalization in line $6$ of Algorithm 2 ensures that the $\lambda$ penalty stays proportional to the penalty for each misclassified point. Our objective (Equation 1) assigns a $\frac{1}{|D|}$ penalty for each misclassification. If we call GOSDT on a smaller subproblem $D(u)$, our penalty per misclassification goes up by a factor of $\frac{|D(u)|}{|D|}$, so we need to scale $\lambda$ to stay proportional.

### v3

@@ 1 Introduction
Decision tree optimization is core to interpretable machine learning (rudin2022interpretable). Simple decision trees present the entire model reasoning process transparently, directly allowing faithful interpretations of the model (arrieta2020explainable). This helps users choose whether to trust the model and to critically examine any perceived flaws.
@@
Figure 1: An illustration of the power of our optimization algorithm. We train 33 decision trees on the Bike dataset, with the aim of predicting bike rentals in Washington DC in a given time period. A greedy tree is fast but suboptimal. An optimal tree is well performing but very slow. Our algorithm strikes the perfect balance, providing well performing trees in a SPLIT second, orders of magnitude faster than optimal approaches seen in literature.
@@
Optimizing the performance of decision trees while preserving their simplicity presents a significant challenge. Traditional greedy methods scale linearly with both dataset size and the number of features (breiman1984classification; quinlan2014c45).
@@
However, these methods tend to yield suboptimal results, lacking general guarantees on either sparsity or accuracy. Recent advances in decision tree algorithms use dynamic programming techniques combined with branch-and-bound strategies, offering solutions that are faster than brute-force approaches and provably optimal (gosdt; dl85; murtree; gosdt_guesses). In fact, murtree and van2024optimal reveal an average gap of 11-22 percentage points between greedy and optimal trees, with murtree showing that some datasets can exhibit gaps as large as $10$ percentage points. These algorithms struggle to scale to datasets with hundreds or thousands of features or to deeper trees. It seems that we should return to greedy methods for larger-scale problems, but this would come at a loss of performance. Ideally, we should leverage greed only when it does not significantly deviate from optimality and use dynamic programming otherwise. Dynamic programming approaches build trees recursively, downward from the root. Problems farther from the root contain fewer samples and produce fewer splits. As we show, greedy splits near the root sacrifice performance, while greedy splits near the leaves produce performance close to the optimal. This suggests that we can tolerate less precision on problems close to leaves than on problems closer to the root – and that full optimization on those problems closer to the leaves yields only marginal returns relative to greedy, since we only have a few splits remaining. This has enormous implications, since the number of candidate trees increases exponentially with increases in depth; using greedy splitting closer to the leaves of the tree massively reduces the search space.
[…]
@@
- We develop a family of decision tree algorithms that scale with the dataset size and number of features comparably to standard greedy algorithms but produce trees that are as accurate and sparse as optimal ones (e.g., gosdt).
- We extend our decision tree algorithms to allow scalable, accurate approximations of the Rashomon set of decision trees (breiman2001statistical; xin2022treefarms).
@@ 2 Related Work
Consistent with recommendations from rudin2022interpretable; costa2023recent, we emphasize sparsity, expressed in terms of the number of leaves, as the primary mechanism for tree interpretability. Sparsity has a strong correlation with user comprehension (Piltaver2016Comprehensible). zhou2018measuring fit a regression model to user-reported interpretability for decision trees, also finding that trees with fewer leaves were more interpretable. They also found that deep, sparse trees were more interpretable than shallow trees with the same sparsity. izza2022tackling provides a way to use a sparse decision tree to provide succinct individual explanations. However, finding deep, sparse trees with existing methods can be computationally infeasible. We bridge this gap – our algorithms are capable of finding sparse trees without constraining them to be shallow.
@@
A long line of work explores greedy algorithms such as CART (breiman1984classification) and C4.5 (quinlan2014c45). These methods first define a heuristic feature quality metric such as the Gini impurity score (breiman1984classification) or the information gain (quinlan2014c45) rather than choosing a global objective function. At every decision node, the feature with the highest quality is chosen as the splitting feature. This process is repeated until a termination criteria is reached. One such criteria often used is the minimum support of each leaf. Trees can then be postprocessed with pruning methods.
@@
Among the many methods for globally optimizing trees, Branch-and-bound approaches with dynamic programming are state of the art for scalability, because they exploit the structure of decision trees (costa2023recent; gosdt; murtree; gosdt_guesses; dl85). While many other methods exist for optimizing trees, such as MIP solvers (bertsimas2017optimal; verwer2019learning), we focus our discussion and comparison of globally optimal decision tree methods on the currently fastest types of approaches – dynamic programming with branch and bound (DPBnB). These approaches search through the space of decision trees while tracking lower and upper bounds of the overall objective at each split to reduce the search space. They can find optimal trees on medium-sized datasets with tens of features and shallow maximum tree depths (maptree; dl85; gosdt; murtree). dl85 uses a DPBnB method with advanced caching techniques to find optimal decision trees, though it does not explicitly optimize for sparsity. In contrast, gosdt; osdt use a DPBnB approach to find a tree that optimizes a weighted combination of empirical risk and sparsity, defined by the number of leaves in the tree. gosdt_guesses further enhances this approach by incorporating smart guessing strategies to construct tighter lower bounds for DPBnB, resulting in computational speedups. murtree extends the work of dl85 by focusing on finding the optimal tree with a hard constraint on the number of permissible nodes, using advanced caching techniques and an optimized depth-2 decision tree solver. quantbnb addresses continuous features by defining lower and upper bounds based on quantiles of feature distributions. However, their method is applicable only to shallow optimal trees with depth $\leq 3$, limiting its utility in scenarios with higher-order feature interactions.
@@
Some older approaches to greedy decision tree optimization consider multiple levels of splits before selecting the best split at a given iteration (Norton1989GeneratingBD). That is, unlike the other greedy approaches, these approaches do not pick the split that optimizes a heuristic immediately. Instead, they pick a split that sets up the best possible heuristic value on the following split.
These approaches still focus on locally optimizing a heuristic measure that is not necessarily aligned with a global objective. By contrast, our method selects splits to directly optimize the sparse misclassification rate of the final tree. We globally optimize the search up to the specified lookahead depth, switching to heuristics only when deciding splits past our lookahead depth. In so doing, our method largely avoids the pathology noted in murthy1995lookahead, who note cases where their own lookahead approach results in a substantially worse tree than one constructed with a standard greedy approach. For our method, it is provably impossible for a fully greedy entropy-based method with the same constraints as our approach to achieve a better training set objective than our approach. (See Theorem A.1)
@@
Several other approaches are compatible with branch and bound techniques. topk seek to bridge the gap between greedy and optimal decision trees by selecting a fixed subset of the top $k$ feature splits for each sub-problem. However, this framework does not explicitly account for sparsity. Further, the method is limited by using a global setting for search precision: the approach considers the same number of candidate splits at each subproblem. As we show in our experiments, there is merit to tailoring the level of search precision to parts of the search space where it is most needed. The Blossom algorithm (demirovic2023blossom) traverses a branch and bound dependency graph structure while using greedy heuristics to guide the search order. Relative to our approach, this algorithm optimizes from the bottom up, starting with greedy splits at each level, then optimizing the splits furthest from the root first. This choice guarantees eventual optimality while giving anytime behavior, but misses out on leveraging the property motivating this work – that greedy splits are most detrimental near the top of the tree. Like the approach of topk, Blossom also does not account for sparsity.
There are a few methods that use probabilistic search techniques to optimize trees. maptree take a Bayesian approach, finding the maximum-a-posteriori tree by optimizing over an AND/OR graph, akin to the graph used in earlier branch-and-bound methods like that of gosdt. Although their method demonstrates strong performance, their experimental results reveal that it is not responsive to sparsity-inducing hyperparameters – accordingly, we found in our experiments that the method struggles to optimize for sparsity.
Recent work by thompson_aistats devises a Monte Carlo Tree Search algorithm using Thompson sampling to enable online, adaptive learning of sparse decision trees. We show that our method achieves superior performance and sparsity on all datasets tested.
@@ 3 Preliminaries
We now briefly discuss sparse greedy and optimal trees. We define $T_{g}(D,d,\lambda)$ to be a decision tree of depth at most $d$ trained greedily on $D$ with sparsity penalty $\lambda$. Intuitively, this sparse greedy algorithm will make a split at a node only when the gain in overall accuracy is greater than $\lambda$. Algorithm 4 in the Appendix illustrates this procedure. Modern methods such as gosdt; gosdt_guesses, on the other hand, find a tree $T$ in the space of decision trees $\mathcal{T}$ that solves the following optimization problem:
@@
$=\min_{T\in\mathcal{T}}\sum_{i=1}^{|D|}\frac{1}{N}\Big({l\big(T(\mathbf{x}_{i}),y_{i}\big)+\lambda S(T)\Big)}\textrm{s.t. depth$(T)$ $\leq d$}$
where $L(T,D,\lambda)$ is the regularized loss of tree $T$ on dataset (or data subset) $D$, $S(T)$ is the number of leaves in $T$, $\ell(T(\mathbf{x}),y)$ is the loss incurred by $T$ in its prediction on $\mathbf{x}$ (for this paper, we set $\ell$ to be the 0-1 loss), and $N$ is the global dataset size. As discussed in Section 2, the fastest contemporary methods solve this problem using a branch-and-bound approach (costa2023recent; gosdt; murtree; gosdt_guesses).
@@
Our work is motivated by the properties of near-optimal decision trees and allows for scalable approximation of that set. xin2022treefarms define the Rashomon set, denoted by $\mathcal{R}(D,\lambda,\epsilon,d)$, as the collection of all trees whose objective is within $\epsilon$ of the minimum value in Equation 1. Formally:
@@
Rashomon sets can be used for a range of downstream tasks (RudinEtAlAmazing2024); one crucial task is the measurement of variable importance over a set of near-optimal models instead of only for a single model (donnelly2023the; fisher2018model). Reliable variable importance measures in this setting rely on minimal feature selection prior to computing the Rashomon set and minimal constraints on the tree’s depth to allow high-order interactions. Our approach can be used to accelerate the computation of a Rashomon set, supporting the feasibility of these approaches.
@@
Given a depth budget $d$, branch and bound with a sparsity penalty (gosdt; gosdt_guesses) finds the optimal loss $\mathcal{L}^{*}(D,d,\lambda)$ that minimizes Equation 1.
@@
Our algorithm will assume feature vectors to be binary, i.e., $\mathbf{x}_{i}\in\{0,1\}^{K}$. Real-world datasets often have features that require discretization to fit our setting. While some methods preserve optimality (e.g., splitting at the mean between unique values in the training set), others such as bucketization (described and proven to be suboptimal in gosdt), binning into quantiles, and feature engineering reduce the search space at the cost of optimality. In our experiments, we use threshold guessing (gosdt_guesses), which sacrifices optimality with respect to a real-valued dataset but maintains theoretical and empirical guarantees relative to a reference decision tree ensemble.
@@ 4 Algorithm Motivation
A key motivating property of SPLIT is that we can find high quality trees even when splitting greedily far from the root of the tree. To support this intuition, we empirically investigate how frequently near optimal trees behave greedily far from the root. To do so, we first generate the Rashomon set of decision trees for various values of sparsity penalty $\lambda$ and Rashomon bound $\epsilon$. Let $T\in\mathcal{R}(D,\lambda,\epsilon,d)$ be a tree in the Rashomon set, and let $n\in T$ be any node in $T$. Then, we compute the fraction of all nodes at a given level $\ell\leq d$ (where level 0 corresponds to the root) that were greedy (by which we mean that the split at this node in the tree is optimal with respect to information gain). This corresponds to the following proportion:
@@
Figure 2: A heatmap of the proportion of splits of trees in the Rashomon set that are greedy, stratified by level, for different ($\lambda,\epsilon$) combinations. Only 44 levels are shown as the $5^{th}$ level corresponds to the leaf. The greyed out regions in the bottom right of a plot represent ($\lambda,\epsilon$) for which the Rashomon set did not contain any trees of that depth. Generally, as we approach the leaves, the proportion of splits appearing in $\epsilon$-optimal trees become increasingly greedy. This is especially noticeable for the Netherlands, Covertype, and COMPAS datasets.
@@ 5.1 SParse Lookahead for Interpretable Trees (SPLIT)
$\begin{cases}\begin{aligned} &\min\Bigg\{\lambda+\frac{|D^{-}|}{N},\lambda+\frac{|D^{+}|}{N},\\ &\min_{f\in\mathcal{F}}\Big\{L\Big(T_{g}\big(D(f),d^{\prime},\lambda\big)\Big)+L\Big(T_{g}\big(D(\bar{f}),d^{\prime},\lambda\big)\Big)\Big\}\Bigg\}\end{aligned}\ \ \\ \hskip 142.26378pt\text{if $d^{\prime}=d-d_{l}$}\\ \begin{aligned} &\min\Bigg\{\lambda+\frac{|D^{-}|}{N},\lambda+\frac{|D^{+}|}{N},\\ &\min_{f\in\mathcal{F}}\Big\{\mathcal{L}\Big(D(f),d^{\prime}-1,\lambda\Big)+\mathcal{L}\Big(D(\bar{f}),d^{\prime}-1,\lambda\Big)\Big\}\Bigg\}\end{aligned}\ \ \\ \hskip 142.26378pt\text{if $d^{\prime}>d-d_{l}$.}\end{cases}$ (4)
Algorithm 1 get_bounds($D$, $_{l}, dd$, $d^{\prime}$, NN) $\to$ lb, ub
0:$D$, $_{l}, dd$, $d^{\prime}$, NN {support, lookahead depth, current search depth, maximum search depth, size of full dataset in GOSDT call}
1:if $d^{\prime}=d_{l}$ then
[…]
@@
4:  $\alpha\leftarrow\frac{1}{N}\sum_{(x,y)\in D}\mathbf{1}[y\neq T_{g}(x)]+\lambda S(T_{g})$
@@
9:  $ub\leftarrow\lambda+\min\Big\{\frac{|D^{-}|}{N},\frac{|D^{+}|}{N}\Big\}$
@@
Algorithm 2 SPLIT($\ell$, D, $\lambda$, $_{l}, dd$, $p$)
[…]
1: ModifiedGOSDT = GOSDT reconfigured to use get_\_bounds (Algorithm 1) whenever it encounters a new subproblem
[…]
@@
Where N is the size of the dataset at the root. We can constrain the search space to include only greedy trees past the lookahead depth by modifying the lower and upper bounds used in branch and bound (see Algorithm 1). In particular, sub-problem nodes initialized at depths up to the lookahead depth are assigned initial lower and upper bounds equivalent to that in GOSDT (gosdt) (see Section 2). At the lookahead depth, however, the lower and upper bounds for a subproblem are fixed to be the loss of a greedy subtree trained on that subproblem. After these bound assignments, our algorithm uses the GOSDT algorithm with these new bounds to solve Equation 5.1 – this is summarized by Lines 11-22 in Algorithm 2. We defer more details of the GOSDT algorithm to Section A.12 in the Appendix.
@@
Once we have solved Equation 5.1, we do not need to use greedy sub-trees past the lookahead depth. We can improve our approach by replacing these subtrees with fully optimal decision trees. Lines 33-99 in Algorithm 2 illustrate this. Thus, the performance of the lookahead tree with the aforementioned greedy subtrees is just an upper bound on the objective of the tree our method ultimately finds.
Note that the renormalization in line $6$ of Algorithm 2 ensures that the $\lambda$ penalty stays proportional to the penalty for each misclassified point. Our objective (Equation 1) assigns a $\frac{1}{N}$ penalty for each misclassification, where $N$ is the size of the full dataset with which GOSDT was called. If the original dataset is $D$, then when we call GOSDT on any descendent subproblem $D(u)$, our penalty per misclassification goes up by a factor of $\frac{|D|}{|D(u)|}$. We need to scale $\lambda$ appropriately to stay proportional to the original dataset $D$.